OpenStudy (anonymous):

The point (17, 0) lies on the line 8x + 11y = 136. Determine if this point is a solution of the system 8x + 11y = 136 0x – y = -8

OpenStudy (anonymous):

somone please help

jimthompson5910 (jim_thompson5910):

Is the second one really 0x – y = -8 ???

OpenStudy (anonymous):

YE

OpenStudy (anonymous):

YES

jimthompson5910 (jim_thompson5910):

0x – y = -8 -y = -8 y = 8 Notice how this is not the same as the y coordinate of (17,0). So (17,0) is not a solution

OpenStudy (anonymous):

THAT ALL I HAVE TO DO YOU MADE IT SEEM SO EASY WOW

jimthompson5910 (jim_thompson5910):

yeah that's why I wanted to check with you whether it was 0x-y = -8 or not

OpenStudy (anonymous):

YOU THINK YOU UNDERTAND THIS QUESTION . Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. -x + 2y = 8 -x + 10y = 48

jimthompson5910 (jim_thompson5910):

change the signs of each term in the first equation to get x-2y = -8 Then add the equations x - 2y = -8 -x + 10y = 48 --------------- 0x + 8y = 40 So we have 8y = 40 Solve 8y = 40 for y